# Differentiable Task Graph Learning: Procedural Activity Representation and Online Mistake Detection from Egocentric Videos

### v1

@@ 5 Limitations
The proposed approach requires the availability of key-step sequences, a common assumption of works addressing other video understanding tasks such as action recognition. While our method is applicable to any fully supervised video understanding dataset, future works should focus on overcoming such limitation and taking advantage of the vast amount of unlabeled video and textual data sets. While the proposed TGT method has shown promising results when trained directly on video features, the investigation of task graph learning in the absence of labeled key-step sequences is beyond the scope of this paper. We noted a reduced ability of our approach to work with noisy action sequences and a tendency to hallucinate pre-conditions, likely due to the limited expressivity of key-steps sequences arising from videos showing the most common ways to perform a procedure.
@@ 7 Acknowledgments
Appendix A Evaluation Measures
@@
Appendix B Implementation Details
@@ B.2 Early Stopping
B.3 Hyperparameters
Table 4 details the hyperparameters employed in the experiments for task graph generation on the CaptainCook4D dataset [26]. During the training of TGT, we utilized a pre-trained EgoVLPv2 [28] to extract text and video embeddings. The temperature value $T$ used in the cross-entropy distinctiveness loss was set to $0.9$ as in [29].
For the downstream task of online mistake detection within the DO model framework, we extended the maximum training epochs to 1200 for Assembly101 [33]. This change was necessary because, even after 1000 epochs, the model continued to exhibit many cycles among its 86 nodes. Extending the number of epochs allows the model additional time to learn and minimize these cycles, which is crucial given the complexity of the graph. In the TGT configuration, we preserved the hyperparameters consistent with those used in the CaptainCook4D settings, except for adjusting the beta values to 1.
You are referred to the code for additional implementation details.
Table 4: List of hyperparameters used in the models training process for task graph generation using CaptainCook4D [26].
@@ B.3 Hyperparameters
$\beta$ 0.005 0.55
@@ B.5 Data Split
B.6 Pairwise ordering and future prediction
@@
We trained our TGT model using video embeddings extracted with a pre-trained EgoVLPv2 [28]. During the training process, if multiple video embeddings are associated with the same key-step across the training sequences, one embedding per key-step is randomly selected. The model is trained for graph generation on the training video and tested for pairwise ordering and future prediction on the test set. For pairwise ordering, we feed our model with the two clips and obtain a $2\times 2$ adjacency matrix. We then select the ordering related to the largest off-diagonal weight. This weight denotes a $A\to B$ edge, suggesting a $B,A$ order ($B$ is a precondition of $A$). For future prediction, we feed the three clips and obtain a $3\times 3$ adjacency matrix. We hence inspect the weights of edges $anchor\to A$ and $anchor\to B$ and choose as future clip, the one related to smallest weight (a small weights indicates that the selected clip is not a precondition).
Considered Data
To evaluate the performance of high quality graph representations, for these experiments we consider the top-5 scenarios of CaptainCook4D, according to graph prediction performance obtained by our TGT model, namely: “Microwave Egg Sandwich”, “Ramen”, ‘Tomato Mozzarella Salad”, “Microwave French Toast”, “Coffee”.
B.7 Graph Post-processing
@@
B.8 Details on Online Mistake Detection
@@
B.9 Qualitative Examples
Figure 5 reports a qualitative example of a prediction.
B.10 Experiments Compute Resources
@@
Figure 5: A qualitative example of a graph predicted by our DO model on the “Microwave Egg Sandwich” scenario of CaptainCook4D [26]. Our method correctly predicts all ground truth edges (green ticks), but also tends to include some additional dependencies such as “top cup with salsa” $\to$ “cut muffin” (red crosses), probably due to the dataset covering the most typical ways of performing the procedure.
Appendix C Societal Impact

### v3

@@ 5 Limitations
The proposed approach requires the availability of key-step sequences, a common assumption of works addressing other video understanding tasks [6, 22, 19, 17, 18]. While our method is applicable to any fully supervised video understanding dataset, future works should focus on overcoming such limitation and taking advantage of the vast amount of unlabeled video and textual data sets. While the proposed TGT method has shown promising results when trained directly on video features, the investigation of task graph learning in the absence of labeled key-step sequences is beyond the scope of this paper. We noted a reduced ability of our approach to work with noisy action sequences and a tendency to hallucinate pre-conditions, likely due to the limited expressivity of key-step sequences arising from videos showing the most common ways to perform a procedure. The performance of our designed system to detect mistakes is influenced by the quality of action recognition (see Figure 4). If the action recognition module fails to detect an action, the method may incorrectly signal a missing pre-condition. Conversely, if an action is falsely detected as performed, the method may fail to signal an actual mistake. Future improvements in online action recognition will enhance the robustness of our method. Furthermore, our approach does not explicitly model “optional” key-steps, which can lead to incorrect error signaling if optional steps are treated as mandatory. This issue could potentially be addressed through the integration of specialized modules capable of detecting optional nodes. Another limitation of task graph representations, both in this work and in prior approaches [3, 44, 39, 20], is their inability to explicitly model repeatable key-steps. Recent advancements such as [18] have introduced a “repeatable” node attribute to task graphs, but this extension is based on manual annotations, and the automatic learning of such attributes from data remains an open problem. Despite this limitation, the proposed error detection model demonstrates an ability to handle cases where key-steps may recur (e.g., spreading peanut butter). At test time, pre-conditions of key-steps are verified via the predicted task graph, even if a key-step has appeared earlier in the sequence. Nevertheless, more effective modeling of repeatable key-steps, especially in contexts where specific repetitions are required (e.g., “cut three slices of bread”), remains an important area for future research. Future work should explore methods for incorporating these requirements into task graph learning frameworks. Our method follows the setup of PREGO [13], which defines the Assembly101-O and EPIC-Tent-O datasets as curated versions of their originals to account for open-set procedural errors such as “order”, “omission”, “correction”, and “repetition” mistakes. These are procedural mistakes, as distinguished from “proficiency errors” described in prior works [18]. The proposed method focuses on procedural mistakes at the abstract level of executed actions, and thus, would not be directly applicable to proficiency error detection. In real-world systems, this limitation could be mitigated by integrating subsystems that specialize in detecting different types of errors. Developing an integrated approach that addresses both procedural and proficiency errors is a promising direction for future research.
Figure 4: To further investigate the effect of noise, we conducted an analysis based on the controlled perturbation of ground truth action sequences, with the aim to simulate noise in the action detection process. At inference, we perturbed each key-step with a probability $\alpha$ (the “perturbation rate”), with three kinds of perturbations: insert (inserting a new key-step with a random action class), delete (deleting a key-step), or replace (randomly changing the class of a key-step). The plots show the trend of the F1 score (Average, Correct, and Mistake) as the perturbation rate increases in the case of Assembly101-O (left) and EPIC-Tent-O (right). Results suggest that the proposed approach can still bring benefits even in the presence of imperfect action detections, with the average F1 score dropping down $10-15$ points with a moderate noise level of $20\%$.
@@ 7 Acknowledgments
Appendix A Task Graph
An example of a task graph is illustrated in Figure 5. A task graph is a Directed Acyclic Graph (DAG) where nodes represent key-steps and directed edges impose a partial order on these steps, indicating the necessary preconditions for each node. For example, the key-step “Mix” has preconditions such as “Add Water”, “Add Milk”, and “Crack Egg”. This formulation of task graphs is not a novel contribution of this paper but was originally introduced in [18].
Appendix B Evaluation Measures
@@
Appendix C Implementation Details
@@ C.2 Early Stopping
C.3 Hyperparameters
Table 4 details the hyperparameters employed in the experiments for task graph generation on the CaptainCook4D dataset [30]. During the training of TGT, we utilized a pre-trained EgoVLPv2 [32] on Ego-Exo4D [18] to extract text and video embeddings. The temperature value $T$ used in the cross-entropy distinctiveness loss was set to $0.9$ as in [33]. The $\beta$ parameter was linearly annealed from an initial value of 1.0 to a final value of 0.05, with updates occurring every 100 epochs. This gradual decrease in $\beta$ mimics the warm-up strategy of [40], enabling smoother optimization early in training and leading to improved convergence as training progresses.
Table 5 details the hyperparameters employed in the experiments for task graph generation on the Assembly101-O and EPIC-Tent-O datasets. For the downstream task of online mistake detection within the DO model framework, we extended the maximum training epochs to 1200, particularly for Assembly101-O. This change was necessary because, even after 1000 epochs, the model continued to exhibit many cycles among its 86 nodes. Extending the number of epochs allows the model additional time to learn and minimize these cycles, which is crucial given the complexity of the graph. In the TGT configuration, we reduced the dropout rate, while the $\beta$ parameter was gradually annealed from an initial value of 1.0 to 0.55 to prevent overfitting.
The reader is referred to the code for additional implementation details.
Table 4: List of hyper-parameters used in the models training process for task graph generation using CaptainCook4D [30].
@@ C.3 Hyperparameters
$\beta$ 0.005 1.0 $\sim$ 0.05
Dropout Rate - 0.25
Table 5: List of hyper-parameters used in the models training process for task graph generation using Assembly101-O and EPIC-Tent-O.
Value
Hyper-parameter DO TGT
Learning Rate 0.1 0.000001
Max Epochs 1200 1200
Optimizer Adam Adam
$\beta$ 0.005 1.0 $\sim$ 0.55
@@ C.5 Data Split
C.6 Pairwise ordering and future prediction
@@
We trained our TGT model using video embeddings extracted with a pre-trained EgoVLPv2 [32] on Ego-Exo4D [18]. During the training process, if multiple video embeddings are associated with the same key-step across the training sequences, one embedding per key-step is randomly selected. The model is trained for task graph generation on the training video and tested for pairwise ordering and future prediction on the test set.
For pairwise ordering, we feed our model with two clips and obtain a $4\times 4$ adjacency matrix, where the nodes represent START, $A$, $B$, END. We establish the order between $A$ and $B$ based on the fulfillment of at least one of the following conditions: (a) if the weight of the edge $A\rightarrow B$ is greater than the weight of the edge $B\rightarrow A$, we conclude that $A$ precedes $B$; (b) by analyzing the sequences $<\text{START},A,B,\text{END}>$ and $<\text{START},B,A,\text{END}>$, we calculate their probabilities using Eq. (4). If $P(<\text{START},A,B,\text{END}>\mid Z)$ is greater than $P(<\text{START},B,A,\text{END}>\mid Z)$, we infer that $A$ precedes $B$; (c) if the weight of the edge END $\rightarrow B$ is greater than that of END $\rightarrow A$, it implies that $B$ is a necessary precondition for concluding the procedure, indicating that $B$ follows $A$, and consequently, $A$ precedes $B$. If none of these conditions hold, we determine that $B$ precedes $A$.
For future prediction, we feed three clips and obtain a $5\times 5$ adjacency matrix, where the nodes represent START, $A$, $anchor$, $B$, and END. We hence inspect the weights of edges $anchor\to A$ and $anchor\to B$ and choose as the future clip, the one related to the smallest weight (a small weight indicates that the selected clip is not a precondition). Another method to determine the future clip is by calculating the probabilities of the sequences $<\text{START},A,anchor,B,\text{END}>$ and $<\text{START},B,anchor,A,\text{END}>$ using Eq. (4). If $P(<\text{START},A,anchor,B,\text{END}>\mid Z)$ is greater than $P(<\text{START},B,anchor,A,\text{END}>\mid Z)$, we infer that the sequence involving $A$ before $B$ is more probable, indicating that $B$ is the future clip for $anchor$. Conversely, if the probability of the second sequence is greater, then $A$ is deemed the future clip for $anchor$.
C.7 Scalability of Task Graph Transformer (TGT)
The Direct Optimization (DO) approach requires a separate training session for each procedure. Task Graph Transformer (TGT) offers more flexibility by allowing different sets of key-step embeddings at each forward pass, ideally enhancing scalability. Leveraging this capability, we trained a single TGT text model for all CaptainCook4D procedures. This was achievable due to TGT’s ability to handle varying embeddings per forward pass, enabling simultaneous optimization across multiple procedures during training. As shown in Table 7, the confidence intervals of both the single and unified models highlight some performance variance. The unified model exhibits lower average precision, recall, and F1 scores compared to the individually trained models, with a larger confidence interval. However, the results suggest that TGT-text models can still generalize across diverse procedures, reducing training complexity while maintaining reasonable performance.
Table 7: Performance comparison between the single TGT-text model trained across all CaptainCook4D procedures and the unified model. The confidence intervals in the single models indicate that the unified method performs comparably to training individual models for each procedure.
Method Precision Recall F1
TGT-text (single) 79.9 ±$\pm 8.8$ 81.9 ±$\pm 6.9$ 80.8 ±$\pm 8.0$
TGT-text (unified) 61.5 ±$\pm 12.0$ 68.2 ±$\pm 10.3$ 64.5 ±$\pm 11.9$
C.8 Graph Post-processing
@@
C.9 Details on Online Mistake Detection
@@
In the case of Assembly101, which includes multiple procedural tasks, we opted to consider a single task graph that summarizes all the procedures, rather than generating individual graphs for each.
C.10 Qualitative Examples
Figures 831 report qualitative examples of prediction using our Direct Optimization (DO) method on the procedures of CaptainCook4D. The task graphs must be read in a bottom-up manner, where the START node (bottom) is at the lowest position and represents the first node with no preconditions, while the END node (up) is the final step of the procedure.
Figure 7 reports a qualitative analysis of the generated task graph for detecting the mistakes on EPIC-Tent-O.
C.11 Experiments Compute Resources
@@
Figure 7: A success (left) and failure (right) case on EPIC-Tent-O. Past key-steps’ colors match nodes’ colors. On the left, the current key-step “Pickup/Open Stakebag” is correctly evaluated as a mistake because the step “Pickup/Place Ventcover” is a precondition of the current key-step, but it is not included among the previous key-steps. On the right, “Pickup/Open Supportbag” is incorrectly evaluated as mistake because the step “Spread Tent” is precondition of the current key-step, but it is not included among the previous key-steps. This is due to the fact that our method wrongly predicted “Spread Tent” as a pre-condition of “Pickup/Open Supportbag”, probably due to the two actions often occurring in this order.
Appendix D Societal Impact
@@
Figure 8: (a) Ground truth task graph and (b) predicted task graph of the scenario Breakfast Burritos.
Figure 9: (a) Ground truth task graph and (b) predicted task graph of the scenario Breakfast Burritos.
Figure 10: (a) Ground truth task graph and (b) predicted task graph of the scenario Cheese Pimiento.
Figure 11: (a) Ground truth task graph and (b) predicted task graph of the scenario Coffee.
Figure 12: (a) Ground truth task graph and (b) predicted task graph of the scenario Cucumber Raita.
Figure 13: (a) Ground truth task graph and (b) predicted task graph of the scenario Dressed Up Meatballs.
Figure 14: (a) Ground truth task graph and (b) predicted task graph of the scenario Broccoli Stir Fry.
Figure 15: (a) Ground truth task graph and (b) predicted task graph of the scenario Caprese Bruschetta.
Figure 16: (a) Ground truth task graph and (b) predicted task graph of the scenario Zoodles.
Figure 17: (a) Ground truth task graph and (b) predicted task graph of the scenario Microwave Mug Pizza.
Figure 18: (a) Ground truth task graph and (b) predicted task graph of the scenario Herb Omelet with Fried Tomatoes.
Figure 19: (a) Ground truth task graph and (b) predicted task graph of the scenario Microwave Egg Sandwich.
Figure 20: (a) Ground truth task graph and (b) predicted task graph of the scenario Microwave French Toast.
Figure 21: (a) Ground truth task graph and (b) predicted task graph of the scenario Mug Cake.
Figure 22: (a) Ground truth task graph and (b) predicted task graph of the scenario Pan Fried Tofu.
Figure 23: (a) Ground truth task graph and (b) predicted task graph of the scenario Pinwheels.
Figure 24: (a) Ground truth task graph and (b) predicted task graph of the scenario Spicy Tuna Avocado Wraps.
Figure 25: (a) Ground truth task graph and (b) predicted task graph of the scenario Spiced Hot Chocolate.
Figure 26: (a) Ground truth task graph and (b) predicted task graph of the scenario Tomato Mozzarella Salad.
Figure 27: (a) Ground truth task graph and (b) predicted task graph of the scenario Salted Mushrooms.
Figure 28: (a) Ground truth task graph and (b) predicted task graph of the scenario Ramen.
Figure 29: (a) Ground truth task graph and (b) predicted task graph of the scenario Butter Corn Cup.
Figure 30: (a) Ground truth task graph and (b) predicted task graph of the scenario Scrambled Eggs.
Figure 31: (a) Ground truth task graph and (b) predicted task graph of the scenario Tomato Chutney.